# Uncertainty-Aware Measurement of Scenario Suite Representativeness for Autonomous Systems

### v1

@@ Abstract
Assuring the trustworthiness and safety of AI systems, e.g., autonomous vehicles (AV), depends critically on the data-related safety properties, e.g., representativeness, completeness, etc., of the datasets used for their training and testing. Among these properties, this paper focuses on representativeness—the extent to which the scenario-based data used for training and testing, reflect the operational conditions that the system is designed to operate safely in, i.e., Operational Design Domain (ODD) or expected to encounter, i.e., Target Operational Domain (TOD). We propose a probabilistic method that quantifies representativeness by comparing the statistical distribution of features encoded by the scenario suites with the corresponding distribution of features representing the TOD, acknowledging that the true TOD distribution is unknown, as it can only be inferred from limited data.
We apply an imprecise Bayesian method to handle limited data and uncertain priors. The imprecise Bayesian formulation produces interval-valued, uncertainty-aware estimates of representativeness, rather than a single value. We present a numerical example comparing the distributions of the scenario suite and the inferred TOD across operational categories—weather, road type, time of day, etc., under dependencies and prior uncertainty. We estimate representativeness locally (between categories) and globally as an interval.
@@ I Introduction
Table I summarises the data-related safety properties defined in PD ISO/PAS 8800:2024 [ISO8800_2024], DATA SAFETY (V3.6) [SCSC_DataSafety_2024], and ISO/IEC 5259-1 [ISO5259_2025], which together specify the key dataset qualities required for AI and autonomous-system safety assurance.
TABLE I: Data-related safety properties from PD ISO/PAS 8800:2024 [ISO8800_2024], DATA SAFETY (V3.6) [SCSC_DataSafety_2024], and ISO/IEC 5259-1 [ISO5259_2025].
@@
PD ISO/ PAS 8800 [ISO8800_2024] Accuracy; Completeness; Correctness / Fidelity; Independence of datasets; Integrity; Representativeness; Temporality; Traceability; Verifiability.
@@
DATA SAFETY [SCSC_DataSafety_2024] Integrity; Completeness; Accuracy; Timeliness / Temporality; Verifiability; Fidelity / Representation; Lifetime; Confidentiality; Availability.
ISO/IEC 5259-1 [ISO5259_2025] Portability; Understandability; Auditability.
@@
While properties such as accuracy, completeness, integrity, and traceability (Table. I) concern local or procedural aspects of data quality, representativeness uniquely addresses the global, distributional relationship between a dataset and the real-world conditions it is intended to model. Thus, among those properties, representativeness—defined as “the distribution of data corresponds to the information in the environment of the phenomenon to be captured; it is free of biases” [ISO8800_2024]—is of particular importance.
In existing literature and standards, the term environment carries related but distinct meanings. The notion of environment ranges from Operational Design Domain (ODD) distributions [langner_statistical_2023] to Target Operational Domain (TOD) specifications [ISO34503_2023].
ISO 34503 [ISO34503_2023] introduces the concept of the Target Operational Domain (TOD) and distinguishes it from the Operational Design Domain (ODD) by defining them as follows: the TOD describes the real-world conditions that an Automated Driving System (ADS) may encounter and must be capable of operating in safely, whereas the ODD specifies the conditions under which the ADS is designed to operate safely. In practice, the TOD typically represents a broader and more variable environment than the ODD; however, the standard explicitly notes that the two domains can be supersets of one another. Consequently, any discrepancy between the ODD and TOD (the ODD-TOD gap) introduces uncertainty about whether the available data adequately represent the operational conditions the system will face once deployed [jeyachandran_OASISS_2025]. In this paper, we assume that the TOD is a superset of the ODD. Therefore, representativeness is defined by comparing the statistical distribution of features encoded in the scenario suites with the corresponding distribution of features representing the TOD [^1] For brevity, we will refer to them as the “scenario suite distribution” and the “TOD distribution” throughout this paper..
@@
In the context of AI safety assurance, and following ISO/PAS 8800 [ISO8800_2024], the term data encompasses all information used to train, test, and validate AI components. In the case of autonomous systems, this data is typically scenario-based, describing structured representations of the situations the system may encounter during operation. An study [jeyachandran_OASISS_2025] interprets this concept operationally by treating structured scenarios as the fundamental data that may be used to train and test AI models (scenario suites). Each scenario captures a combination of operational variables such as weather, road type, illumination, etc. Assessing AI system safety involves evaluating data-related safety properties (Table I), which together determine the adequacy of the dataset. Among the key adequacy properties, this paper focuses on representativeness.
@@
The central contribution of this work is therefore the estimation of the TOD joint distribution using a Bayesian framework (statistical inference, shown on the l.h.s of Fig. 1) well-suited for reasoning under data sparsity and uncertainty. Furthermore, we employ imprecise probability to represent uncertainty in the prior distributions used in the Bayesian inference. In general, our work formalizes representativeness as a distributional discrepancy between the empirical scenario suite and the “unknown” TOD distribution, and by deriving a Bayesian framework to quantify it together with its epistemic uncertainty (Fig. 1). Consequently, The resulting discrepancy metric is expressed as an uncertainty-aware estimate interval.
@@
The remainder of this paper is organized as follows. Section II reviews related work on dataset representativeness in AI safety assurance. Section III presents the proposed metric for quantifying representativeness. A numerical example illustrating its application is provided in Sec. IV. Section. V discusses the results, and finally, Sec. VI concludes the paper.
@@ II Related Work
This section reviews existing approaches to representativeness evaluation and imprecise probability to position our probabilistic framework within this landscape.
@@ II-A Comparison with Existing Methods on Representativeness
Recent work in automated driving has recognized that the representativeness of test scenarios is a key determinant of the credibility of safety evidence. Langner et al. [langner_statistical_2023] formalize representativeness as a statistical distance between two empirical distributions: the observed open-road test data and the target operational design domain (ODD) reference statistics.
De Gelder et al. [de_scenario_2022] apply this idea to the comparison between generated scenario data and real-world driving data. They use a combination of singular value decomposition (SVD) and kernel density estimation (KDE) to model the joint distribution of driving parameters, and evaluate how closely the generated scenarios approximate the empirical distribution of real-world data.
Hiller et al. [hiller_creating_2022] shift focus from the metric itself to the process of representativeness assessment. Their framework compares distributions of recorded driving data against reference statistics derived from heterogeneous sources such as OpenStreetMap, meteorological services, traffic counters, and accident databases.
In [jeyachandran_OASISS_2025], representativeness is defined as the similarity between the empirical distribution of operational conditions in the scenario suite and the corresponding distribution in the TOD. Both are modeled as categorical probability distributions derived from ISO-standardized operational factors.
At the assurance level, the JRC report on Trustworthy Autonomous Vehicles [fernandez_trustworthy_2021] identifies representativeness and data coverage as measurable properties required for trustworthy validation and testing. While this establishes representativeness as a normative requirement, it does not specify how to quantify it or account for uncertainty.
The above methods evaluate representativeness as a distance between distributions of two empirical datasets: generated versus real-world [de_scenario_2022], test versus reference ODD data [langner_statistical_2023], or scenario suites versus TOD [jeyachandran_OASISS_2025]. The real-world dataset is treated as a reliable surrogate for the true operational domain, assuming that sufficient data exist to capture its variability. In practice, however, this assumption may not hold when operational data are sparse or partially observed. Some studies, such as Gelder et al [de_scenario_2022], highlight this issue and quantify how sampling bias in real-world testing affects the validity of validation results using frequency-based divergence measures. Our probabilistic framework (Bayesian inference) addresses this by comparing the empirical distribution of the scenario suite with an estimated approximation of the true, but unknown, TOD distribution. Instead of assuming that the TOD can be directly estimated from abundant real-world data, we treat its distribution as uncertain and propagate epistemic uncertainty into the representativeness metrics. Thus, our work extends deterministic distribution-matching approaches into a robust, uncertainty-aware framework.
@@ II-B Imprecise Probability
While classical Bayesian methods provide a coherent mechanism for uncertainty propagation, they depend on the specification of a single prior distribution, which can be challenging to justify under sparse data. Imprecise probability frameworks [augustin2014introduction, troffaes2007decision, utkin_imprecise_2018, imprecision_luck_walter_2009, zhao_interval_2020] relax this assumption by representing uncertainty using set of distributions (credal sets) instead of a single one (Table. II). Estimating such set of posterior distributions poses significant computational challenges. This is because there is generally no monotonic relationship between the prior and the posterior. However, when the prior belongs to the exponential family, a reparameterization can be applied—expressing the prior in terms of its strength and mean parameters—according to the iLUCK model proposed by [imprecision_luck_walter_2009]. In the context of our categorical data, the iLUCK model [imprecision_luck_walter_2009] allows the specification of bounds on these parameters. This formulation preserves conjugacy while enabling monotonic propagation of prior imprecision into posterior intervals. This monotonicity property allows us to compute the posterior bounds efficiently by evaluating only the corner cases of the prior parameter space, rather than calculating the posterior for every possible prior configuration.
@@
III A metric for quantifying the representativeness
Given the definition of TOD and tailored definition of data in this study, it is conceived as a multidimensional space that describes the range of real-world operating conditions under which an AI system is expected to function safely. Each dimension of this space corresponds to an operational variable, such as weather, road type, illumination, traffic density, vehicle speed etc (Table. III).
Because many of these operational variables are continuous in nature—such as illumination intensity, temperature, or speed—it is often easier and more practical to represent the TOD in a finite, analyzable form through discretization (see Table. III). In practice, standards and safety guidelines such as ISO/PAS 8800:2024 [ISO8800_2024] and DATA SAFETY (V3.6) [SCSC_DataSafety_2024] establish the requirement for a well-defined coverage of the operational domain but leave the specific discretization boundaries to be determined by domain-specific standards or, in their absence, through expert judgment.
@@ III-A Estimating the Posterior Distribution Envelopes of TOD
Theorem 1 formalizes how to estimate the lower and upper bounds on the posterior distribution of TOD under imprecise priors.
@@
Let the Target Operational Domain (TOD) be represented by a categorical random variable $Z\in\{1,\ldots,K\}$ with unknown probability vector $\boldsymbol{\theta}_{\text{TOD}}=(\theta_{1},\ldots,\theta_{K})\in\Delta_{K-1}$ where, $\Delta_{K-1}$ is the probability simplex as follows:
@@
Given observed category counts $\mathbf{k}=(k_{1},\ldots,k_{K})$ from $n$ i.i.d. samples, the likelihood is multinomial:
@@
The joint posterior forms a family of Dirichlet distributions over the simplex, whose lower and upper envelopes are attained at the extreme points of $\mathcal{Y}_{\text{TOD}}$ and $\underline{n},\overline{n}]$. Marginally, each $\theta_{k}$ follows a corresponding family of Beta distributions, with posterior bounds obtained in the same way.
See Appendix A for a proof outline.
@@ III-B Representativeness Metric under Uncertainty
Having established a method to infer the posterior distribution of the TOD in Theorem 1, we now address how to quantify the representativeness of the scenario suite by measuring the discrepancy between two distributions: the empirical scenario suite distribution $\hat{\boldsymbol{\pi}}_{\mathrm{suite}}$ and the posterior TOD distribution $\boldsymbol{\theta}_{\mathrm{TOD}}|{\bf k}$.
A representativeness metric must capture the overall alignment between these two categorical distributions defined over the same $K$ joint categories. We employ two complementary measures that are well-established in information theory and probability theory:
@@ III-B3 Propagating Epistemic Uncertainty
These bounds illustrate epistemic uncertainty in representativeness, indicating how sensitive the assessment is to prior beliefs about the TOD. A wide ones suggest the need for more operational data to refine the estimate.
@@ IV Numerical Example
This section presents numerical examples illustrating the proposed method.
@@ IV-A Problem Setting: Common Cases
The TOD and the scenario suite share the same variable set, discretization, and categorical space. Both are defined over five operational factors: $W,R,D,T,S,$ where
  • Weather $W\in\{\text{Clear},\text{Adverse}\}$,
  • Road $R\in\{\text{Highway},\text{Urban}\}$,
  • Time $D\in\{\text{Day},\text{Night}\}$,
  • Traffic $T\in\{\text{Light},\text{Heavy}\}$,
  • Speed $S\in\{\leq 50,\;>50\}$.
Each combination $(w,r,d,t,s)$ defines one of $K=2^{5}=32$ categories (Table IV), which serve as a common comparison basis for both the TOD and the scenario suite.
@@ 31 Adverse Urban Night Heavy >50>50
To ensure realism and consistency, dependencies among these factors are explicitly incorporated in both models through a shared Bayesian-network structure $(R\!\to\!W,\;D\!\to\!T,\;R\!\to\!S)$. This structure captures operationally relevant relationships such as weather conditions depending on road type, traffic patterns varying with time of day, and speed distributions differing across road environments. This structure is applied consistently to both the scenario suites and the Target Operational Domain (TOD). It captures realistic relationships among scenario variables: road type influences both weather exposure and typical speed, while time of day affects traffic conditions. The root marginal distributions are specified as
@@
The corresponding conditional probability tables (CPTs) are summarized in Table. V.

### v2

@@ Abstract
Assuring the trustworthiness and safety of AI systems, e.g., autonomous vehicles (AV), depends critically on the data-related safety properties, e.g., representativeness and completeness of the datasets used for their training and testing. Among these properties, this paper focuses on representativeness—the extent to which the scenario-based data used for training and testing, reflect the operational conditions that the system is designed to operate safely in, i.e., Operational Design Domain (ODD) or expected to encounter, i.e., Target Operational Domain (TOD). We propose a probabilistic method that quantifies representativeness by comparing the statistical distribution of features encoded by the scenario suites with the corresponding distribution of features representing the TOD, acknowledging that the true TOD distribution is unknown, as it can only be inferred from limited data. We apply an imprecise Bayesian method to handle limited data and uncertain priors. The imprecise Bayesian formulation produces interval-valued, uncertainty-aware estimates of representativeness, rather than a single value. We present numerical examples comparing the distributions of the scenario suite and the inferred TOD across operational categories (e.g., weather and road type) under dependencies and prior uncertainty. We estimate representativeness locally (between categories) and globally as an interval.
@@ I Introduction
Table I summarises data-related safety properties defined in PD ISO/PAS 8800:2024 [9], DATA SAFETY (V3.6) [4], and ISO/IEC 5259-1 [2], which together specify key dataset qualities for AI and autonomous-system safety assurance.
TABLE I: Data-related safety properties from PD ISO/PAS 8800:2024 [9], DATA SAFETY (V3.6) [4], and ISO/IEC 5259-1 [2].
@@
PD ISO/ PAS 8800 [9] Accuracy; Completeness; Correctness / Fidelity; Independence of datasets; Integrity; Representativeness; Temporality; Traceability; Verifiability.
@@
DATA SAFETY [4] Integrity; Completeness; Accuracy; Timeliness / Temporality; Verifiability; Fidelity / Representation; Lifetime; Confidentiality; Availability.
ISO/IEC 5259-1 [2] Portability; Understandability; Auditability.
@@
While properties such as accuracy, completeness, integrity, and traceability (Table I) concern local or procedural aspects of data quality, representativeness uniquely addresses the global, distributional relationship between a dataset and the real-world conditions it is intended to model. Thus, among those properties, representativeness—defined as “the distribution of data corresponds to the information in the environment of the phenomenon to be captured; it is free of biases” [9]—is of particular importance.
In existing literature and standards, the term environment carries related but distinct meanings. The notion of environment ranges from Operational Design Domain (ODD) distributions [11] to Target Operational Domain (TOD) specifications [8]. ISO 34503 [8] distinguishes between ODD, which defines the conditions under which an ADS is designed to operate safely, and TOD, which describes the real-world conditions the ADS may encounter and must handle safely. The standard allows different relationships between these domains, including partial overlap or one being broader than the other [^1] In practice, the TOD is often broader than the ODD.. We restrict representativeness assessment to the intersection $\mathrm{TOD}\cap\mathrm{ODD}$, considering only ODD-compliant scenarios. Representativeness is therefore evaluated by comparing the statistical distribution of features (i.e., scenario attributes such as road, weather, and traffic conditions) in the scenario suites with the corresponding distribution of features representing this restricted TOD [^2] For brevity, we will refer to them as the “scenario suite distribution” and the “TOD distribution”..
@@
In the context of AI safety assurance, and following ISO/PAS 8800 [9], the term data encompasses all information used to train, test, and validate AI components. In the case of autonomous systems, this data is typically scenario-based, describing structured representations of the situations the system may encounter during operation. A study [10] interprets this concept operationally by treating structured scenarios as the fundamental data that may be used to train and test AI models (scenario suites). Each scenario captures a combination of operational variables such as weather, road type, illumination, etc. Assessing AI system safety involves evaluating data-related safety properties (Table I), which together determine the adequacy of the dataset. Among the key adequacy properties, this paper focuses on representativeness.
@@
The central contribution of this work is therefore the estimation of the TOD joint distribution using a Bayesian framework (statistical inference, shown on the l.h.s of Fig. 1) well-suited for reasoning under data sparsity and uncertainty. Furthermore, we employ imprecise probability to represent uncertainty in the prior distribution used in the Bayesian inference. In general, our work formalizes representativeness as a distributional discrepancy between the empirical scenario suite and the “unknown” TOD distribution, and by deriving a Bayesian framework to quantify it together with its epistemic uncertainty (Fig. 1). Consequently, the resulting discrepancy metric is expressed as an uncertainty-aware estimate interval.
@@
The remainder of this paper is organized as follows. Section II reviews related work on dataset representativeness in AI safety assurance. Section III presents the proposed metric for quantifying representativeness. A numerical example illustrating its application is provided in Sec. IV. Section V discusses the results, and finally, Sec. VI concludes the paper.
@@ II Related Work
This section reviews existing approaches to representativeness evaluation and imprecise probability to position the proposed framework.
@@ II-A Comparison with Existing Methods on Representativeness
Recent work in automated driving has recognized that the representativeness of test scenarios is a key determinant of the credibility of safety evidence. Langner et al. [11] formalize representativeness as a statistical distance between two empirical distributions: the observed open-road test data and the target operational design domain (ODD) reference statistics. De Gelder et al. [5] apply this idea to the comparison between generated scenario data and real-world driving data. They use a combination of singular value decomposition (SVD) and kernel density estimation (KDE) to model the joint distribution of driving parameters, and evaluate how closely the generated scenarios approximate the empirical distribution of real-world data. Hiller et al. [7] shift focus from the metric itself to the process of representativeness assessment. Their framework compares distributions of recorded driving data against reference statistics derived from heterogeneous sources such as OpenStreetMap, meteorological services, traffic counters, and accident databases. In [10], representativeness is defined as the similarity between the empirical distribution of operational conditions in the scenario suite and the corresponding distribution in the TOD. Both are modeled as categorical probability distributions derived from ISO-standardized operational factors. At the assurance level, the JRC report on Trustworthy Autonomous Vehicles [6] identifies representativeness and data coverage as measurable properties required for trustworthy validation and testing. While this establishes representativeness as a normative requirement, it does not specify how to quantify it or account for uncertainty.
The above methods evaluate representativeness as a distance between distributions of two empirical datasets: generated versus real-world [5], test versus reference ODD data [11], or scenario suites versus TOD [10]. The real-world dataset is treated as a reliable surrogate for the true operational domain, assuming that sufficient data exist to capture its variability. In practice, however, this assumption may not hold when real world data are sparse or partially observed. Some studies, such as Gelder et al [5], highlight this issue and quantify how sampling bias in real-world testing affects the validity of validation results using frequency-based divergence measures. Our probabilistic framework (Bayesian inference) addresses this by comparing the empirical distribution of the scenario suite with an estimated approximation of the true, but unknown, TOD distribution. Instead of assuming that the TOD can be directly estimated from abundant real-world data, we treat its distribution as uncertain and propagate epistemic uncertainty into the representativeness metrics. Thus, our work extends deterministic distribution-matching approaches into a robust, uncertainty-aware framework.
@@ II-B Imprecise Probability
While classical Bayesian methods provide a coherent mechanism for uncertainty propagation, they depend on the specification of a single prior distribution, which can be challenging to justify under sparse data. Imprecise probability frameworks [1] relax this assumption by representing uncertainty using set of distributions (credal sets) instead of a single one (Table II). Estimating such set of posterior distributions poses significant computational challenges. This is because there is generally no monotonic relationship between the prior and the posterior. However, when the prior belongs to the exponential family, a reparameterization can be applied—expressing the prior in terms of its strength and mean parameters—according to the iLUCK model proposed by [12] and applied in different studies ([13, 14]). In the context of our categorical data, the iLUCK model [12] allows the specification of bounds on these parameters. This formulation preserves conjugacy while enabling monotonic propagation of prior imprecision into posterior intervals. This monotonicity property allows us to compute the posterior bounds by evaluating only the corner cases of the prior parameter space, rather than calculating the posterior for every possible prior configuration.
@@
III A representativeness metric
Given the definition of TOD and tailored definition of data in this study, it is conceived as a multidimensional space that describes the range of real-world operating conditions under which an AI system is expected to function safely. Each dimension of this space corresponds to an operational variable, such as weather, road type, illumination, etc., (Table III).
Because many of these operational variables are continuous in nature—such as illumination intensity, temperature, or speed—it is often easier and more practical to represent the TOD in a finite, analyzable form through discretization (see Table III). In practice, standards and safety guidelines such as ISO/PAS 8800:2024 [9] and DATA SAFETY (V3.6) [4] establish the requirement for a well-defined coverage of the operational domain but leave the specific discretization boundaries to be determined by domain-specific standards or, in their absence, through expert judgment.
@@ III-A Estimating the Posterior Distribution Envelopes of TOD
Theorem 1 formalizes the computation of lower and upper bounds on the TOD posterior under imprecise priors.
@@
Let the Target Operational Domain (TOD) be represented by a categorical random variable $Z\in\{1,\ldots,K\}$ with unknown probability vector $\boldsymbol{\theta}_{\text{TOD}}=(\theta_{1},\ldots,\theta_{K})\in\Delta_{K-1}$ where, $\Delta_{K-1}$ is all valid probability distributions over $K$ categories:
@@
Given observed category counts $\mathbf{k}=(k_{1},\ldots,k_{K})$ from $n$ i.i.d.333i.i.d. means that each scenario is treated as an independent draw from the TOD distribution, so the outcome of one scenario does not affect another. samples, the likelihood is multinomial:
@@
The joint posterior forms a family of Dirichlet distributions over the simplex, whose lower and upper envelopes are attained at the extreme points of $\mathcal{Y}_{\text{TOD}}$ and $\underline{n},\overline{n}]$. Marginally, each $\theta_{k}$ follows a corresponding family of Beta [^4] This is the mathematical property of the Dirichlet. distributions, with posterior bounds obtained in the same way.
See Appendix A for the outline of proof.
@@ III-B Representativeness Metric under Uncertainty
Given the posterior TOD distribution from Theorem 1, representativeness is quantified as the discrepancy between $\hat{\boldsymbol{\pi}}$ (empirical scenario suite distribution) and $\boldsymbol{\theta}_{\mathrm{TOD}}\mid\mathbf{k}$. Representativeness is quantified by the alignment between these two categorical distributions over the same $K$ joint categories, using two complementary measures from information and probability theory:
@@ III-B3 Propagating Epistemic Uncertainty
These bounds illustrate epistemic uncertainty in representativeness, indicating how sensitive the assessment is to prior beliefs about the TOD. Wide intervals suggest the need for more operational data to refine the estimate.
@@ IV Numerical Example
The numerical example presented in this section is intentionally simplified and uses a small number of abstract categorical variables to illustrate the proposed method in a transparent and reproducible manner.
@@ IV-A Problem Setting: Common Cases
The TOD and the scenario suite share the same variable set, discretization, and categorical space. Both are defined over five operational factors: $W,R,D,T,S,$ where, Weather $W\in\{\text{Clear},\text{Adverse}\}$, Road $R\in\{\text{Highway},\text{Urban}\}$, Time $D\in\{\text{Day},\text{Night}\}$, Traffic $T\in\{\text{Light},\text{Heavy}\}$, and Speed $S\in\{\leq 50,\;>50\}$. Each combination $(w,r,d,t,s)$ defines one of $K=2^{5}=32$ categories (Table IV), which serve as a common comparison basis for both the TOD and the scenario suite.
@@ 31 Adverse Urban Night Heavy >50>50
To ensure realism and consistency, dependencies among these factors are explicitly incorporated in both models through a shared Bayesian-network structure $(R\!\to\!W,\;D\!\to\!T,\;R\!\to\!S)$. This structure captures operationally relevant relationships such as weather conditions depending on road type, traffic patterns varying with time of day, and speed distributions differing across road environments. This structure is applied consistently to both the scenario suites and TOD. It captures realistic relationships among scenario variables: road type influences both weather exposure and typical speed, while time of day affects traffic conditions. The root marginal distributions are specified as
@@
The corresponding conditional probability tables (CPTs) are summarized in Table V.